Racecar A is traveling on a straight portion of the track while racecar B is traveling on a circular portion of the track. ThRacecar A is traveling on a straight portion of the track while racecar B is traveling on a circular portion of the track. ThPlease check this!
Answer #1

Answer Given, VA=250 km/h = 250 x 1000 m/s 3600 = 69.44 m/s = 58.33 m/s(7) V8 = 210 kmh = 210 X 1000 3600 m/s JA =-69.44 i JB» abla =122 K22.544)+(-2.694) m/s2 AB/A = 22.704 Aug Ortañ :2.694 =-6.815° 22.544 => O = 6.815° Aug Good Luck Please like.So your answers are absolutely correct. Please rate.